OpenStudy (anonymous):

what is the slope of the line passing throw the points (0,-5) & (4,2) ?

OpenStudy (anonymous):

Also this one. http://www.mathwarehouse.com/algebra/linear_equation/slope-of-a-line.php

OpenStudy (aaliyahfairgood):

(2+5)/(4-0) =7/4 using slope formula

OpenStudy (hockeychick23):

slope= y2-y1/x2-x1